OpenStudy (anonymous):

The prime factorization of 76 is _____. a) 2 • 38 b) 2^2• 19 c) 2^3• 7 d) 76

hartnn (hartnn):

first divide it by lowest prime, 2 , if possible so, 76/2 = 38 can it be still divided ? yes, so divide it again 38/2 = 19 can it be still divided ? no so u have 76 = 2*2* 19 = 2^2*19

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

Yeah, thanks.

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]